OpenStudy (anonymous):

Solve the following system of equations by linear combination: 4x + 3y = 12 y = x – 10 (–4, 6) (6, –4) (–7, 3) (3, –7)

OpenStudy (anonymous):

4x + 3y = 12 y = x - 10 ---> -x + y = -10 --> x - y = 10 so we have : 4x + 3y = 12 x - y = 10 --->(-4) ------------ 4x + 3y = 12 -4x + 4y = -40 (result of multiplying by -4) -----------add 7y = - 28 y = -4 now sub -4 in for y in either of the equations y = x - 10 -4 = x - 10 -4 + 10 = x 6 = x check... 4x + 3y = 12 4(6) + 3(-4) = 12 24 - 12 = 12 12 = 12 (correct) x = 6 and y = -4 or (6,-4)